5+2y/3=25/12+5y+3/4

Answer by Boreal(15235)

5+(2y/3)=(25/12)+5y+(3/4). This is how I interpret the above.
multiply everything by 12
60+8y=25+60y+9
26=52y
y=(1/2)
5 1/3=(25/12)+(5/2)+(3/4), which is true.
ONLY IF I WRITE THE PROBLEM THIS WAY, with parentheses used appropriately so as to state what operations should be ordered how. 5+2y/3 is NOT (5+2y)/3. Look it up.
(5+2y)/3=(25/12)+(5y+3)/4
THEN MULTIPLY BY 12
20+8Y=25+15Y+9
THEN y=-2
please study order of operations and use parentheses appropriately. And please do not use statements that belittle those who make an effort to try to help you. My statement "This is how I interpret the above." If that interpretation is in error, then you should have re-submitted the problem instead of telling me the answer and letting me figure out how you should have written the problem in the first place.